The embers in a fireplace are at 812 ºC and have an exposed area of 0.17 m2 and emissivity of 0.931. The surrounding room has a temperature of 16.19ºC. If 60% of the radiant energy enters the room, what is the net rate of radiant heat transfer in kilowatts?


1
Expert's answer
2021-03-23T11:08:57-0400
"P=\eta \sigma eA(T_2^4-T_1^4)\\P=(0.931)(0.6)(0.17(5.67\cdot10^{-8})(812^4-16.19^4)\\P=2340\ W=2.34\ kW"